OpenStudy (anonymous):

HELP!GIVING OUT MEDALS!! Which o the following statements is true about a collision between two cars? A.The total momentum after the collision is the same as the total momentum before the collision B.The total momentum after the collision is greater than the total momentum after the collision is greater than the total momentum before the collision C.The total momentum after the collision is less than the total momentum before the collision D.The total momentum after the collision is zero

OpenStudy (anonymous):

It is A, conservation of momentum It is only A if it is the total linear momentum

OpenStudy (anonymous):

If this was cue balls bouncing off each other, the answer would be A, because they would rebound with the same momentum they had a collision with But since this is cars...and they don't bounce after collisions and have zero velocity then I think the answer for this is D) The total momentum after the collision is zero

OpenStudy (anonymous):

Yes yes, if it is a elastic car collision, then it is D

OpenStudy (anonymous):

This time I'm going with C... C.The total momentum after the collision is less than the total momentum before the collision An elastic collision is one in which no kinetic energy is lost in the collision...like Newton's cradle balls smacking each other. And an inelastic collision is one where KE is lost in the collision. Two cars colliding would lose kinetic energy because of the metal crumpling and absorbing shock and converting that energy to heat. But the cars probably wouldn't come to a complete stop in the collision, so some of the energy is saved as momentum Look at these equations , m1 is car one and m2 is car two : KEf = kinetic energy final, KEi , is kinetic energy initial Ratio of kinetic energies before and after collision: KEf / KEi = m1 / (m1 + m2) Fraction of KE lost in the collision: KEi - KEf / KEi = m / (m1 + m2) So the momentum going into the collision is the same minus the KE lost in the collision, which absorbs some of that momentum.

OpenStudy (anonymous):

.-. didn't mean elastic in that way, I just made a new 'term' where the KE would be neutralized. Well, I guess C answers it in a more general state, since if momentum is 0 then it is of course less than the beginning momentum

OpenStudy (anonymous):

it doesnt matter whether the collision is elastic or inelastic,, momentum is ALWAYS conserved in any collision,,,,provided EXTERNAL forces dont act on the system ,,

OpenStudy (anonymous):

answer must be A

OpenStudy (anonymous):

@LastDayWork

OpenStudy (anonymous):

loss of KE doesnt mean loss of momentum,,,,,

OpenStudy (lastdaywork):

Yea A is the correct answer..unless someone is considering the friction on tires :P

OpenStudy (lastdaywork):

Linear momentum remains conserved even in perfectly inelastic collisions..

OpenStudy (anonymous):

Ah yeah, I forgot one little key thing, yes, my first answer was correct, it is A

OpenStudy (anonymous):

I was rereading the difference between KE and conservation of momentum....and you guys are right..it is answer A) This goes against everything common sense, meaning KE = 1/2mv^2 and momentum = m*v , so it appears with mass and velocity being the only two inputs into what KE and momentum are, that it would make sense if you have less KE after a collision, you would also have less momentum. But you don't, momentum is conserved regardless of the collision type since momentum is considered a vector and by not squaring velocity like you do when calculating KE, that keeps momentum a linear component and not exponential like KE would be. In other words, you can double velocity on KE and KE increases by a factor of four, whereas doubling velocity for momentum only doubles the momentum.

OpenStudy (anonymous):

yeah ,,u are right man

OpenStudy (anonymous):

its quite intuistic to believe that if KE reduces, Momentum also reduces ,, because velocity reduced,,,,,,

OpenStudy (anonymous):

but ,, though velocity reduced......the reduced velocity is with a more MASSIVE BODY ,, so MV remains constant compensating it

OpenStudy (anonymous):

in simple words FORCE= d(P)/dt = constant (means P doesnt change) ONLY when F is a constant ,,,,,,i.e when no force acts on the system , momentum remains intact ,,conserved

OpenStudy (lastdaywork):

"This goes against everything common sense, meaning KE = 1/2mv^2 and momentum = m*v , so it appears with mass and velocity being the only two inputs into what KE and momentum are, that it would make sense if you have less KE after a collision, you would also have less momentum." There's a subtle difference (as you already pointed out), while calculating total KE of the system we add the magnitude of KE of individual parts ; whereas total momentum may be (and generally is) less than the summation of magnitudes of individual parts. In other words, the fact that magnitude of individual velocities decreases after collision DOESN'T imply that the magnitude of their (weighted) difference reduces too.